OpenStudy (jackiem272):

If dy/dx = 7x^2/y^3 and y(3)=2, find an equation for y in terms of x

zepdrix (zepdrix):

\[\Large\bf\sf \frac{dy}{dx}=\frac{7x^2}{y^3},\qquad\qquad y(3)=2\]So this appears to be separable. Do you understand how to get all of the y stuff on one side and all of the x stuff on the other? It will involve moving the differentials (namely dx) as well as the other stuff.

OpenStudy (jackiem272):

Is it y^3dy=7x^2dx? I don't know what to do after this

OpenStudy (jackiem272):

then you get the antiderivative?

zepdrix (zepdrix):

Ok looks good so far. Integrating each side,\[\Large\bf\sf \int\limits y^3\;dy\quad=\quad \int\limits 7x^2\; dx\]

zepdrix (zepdrix):

Yes. Normal power rule on each side.

zepdrix (zepdrix):

Integration Power Rule I mean*

OpenStudy (jackiem272):

y^4/4=7x^3/3?

zepdrix (zepdrix):

Ok good, we have a +c somewhere also right?

OpenStudy (jackiem272):

yes, on the right side?

zepdrix (zepdrix):

Yes, wherever you want it :) Let's put it on the x side.

zepdrix (zepdrix):

Let's also multiply both sides by 4, get the y^4 alone.\[\Large\bf\sf y^4\quad=\quad \frac{28}{3}x^3+c\]

OpenStudy (jackiem272):

can it also be written as y^4= 28x^3/3+C

zepdrix (zepdrix):

Hehe yer too fast for me XD

zepdrix (zepdrix):

Do you understand why it's c and not 4c?

OpenStudy (jackiem272):

And after that I am lost! C is a constant?

zepdrix (zepdrix):

I mean because we multiplied both sides by 4.

zepdrix (zepdrix):

Yes it's a constant, but it's the fact that it's an arbitrary value. It represents a family of solutions. It could be any number. So 4 times any number is still any number. We just absorb the 4 into the c. If you already have a good understanding of that, then keep moving :D

zepdrix (zepdrix):

We `could` solve for y at this point. But I think we're in a good position to use our `initial data` to solve for c.

OpenStudy (jackiem272):

Oh yeah, I've learned that. I don't know what to do after that

OpenStudy (jackiem272):

What do we do with the y(3)=2?

zepdrix (zepdrix):

\[\Large\bf\sf y^4\quad=\quad \frac{28}{3}x^3+\mathcal C\] We plug it in.\[\Large\bf\sf y(3)=2,\qquad\qquad\to\qquad\qquad (2)^4\quad=\quad \frac{28}{3}(3)^3+\mathcal C\]Does that make sense how I plugged the x and y value in?

zepdrix (zepdrix):

Make sure you're comfortable with function notation. \(\Large\bf\sf y(3)=2 \quad\text{means}\qquad\qquad y=2\quad\text{when}\quad x=3\)

OpenStudy (jackiem272):

WOW. I can't believe I forgot about that

zepdrix (zepdrix):

heh :o

OpenStudy (jackiem272):

We have to find C, right? C=-236

zepdrix (zepdrix):

Mmm ya that looks right!

zepdrix (zepdrix):

\[\Large\bf\sf y^4\quad=\quad \frac{28}{3}x^3-236\]

zepdrix (zepdrix):

So how should we finish this up? How do we write it terms of y=?

OpenStudy (jackiem272):

Do we have to write it in terms of y=? We cannot leave it like that?

zepdrix (zepdrix):

`find an equation for y in terms of x` Ya we want it written as y= if it's possible.

OpenStudy (jackiem272):

Agh. do you mind helping me out?

zepdrix (zepdrix):

The inverse of the powers is going to be the root operation. If we had y^2, we would apply the square root (2nd root) to isolate y. The square and root would undo one another, leaving us with y.

zepdrix (zepdrix):

But we have a y^4 right? :o Hmm are there any other root operations we can use?

OpenStudy (jackiem272):

We can't use 4th root?

zepdrix (zepdrix):

Yesss good good.

OpenStudy (jackiem272):

and can we just leave it as y=\[\sqrt[4]{(28x^3/3)-236}\]

zepdrix (zepdrix):

Yes, good. It may not look pretty, but it's correct! :)

OpenStudy (jackiem272):

That's all? Thank you soooooo much! This really helped a lot. The teacher taught this in like 10 minutes haha!

zepdrix (zepdrix):

haha XD np